# package/before/amsmath

### 14 Dec 2023 v1

@@ 3.3.1 Least squares formulation
$(R_{U})^{n}_{m}=(U_{k})^{n+1}_{m}-((\hat{U}_{k})^{n}_{m}-\left(D_{t}(\hat{U}_{k})^{n}_{m}+(\mathcal{L}_{k}\hat{U}_{k})^{n}_{m}\right)\Delta t+\sum\limits_{i=1}^{d}\sigma_{i}D_{S_{i}}(\hat{U}_{k})^{n}_{m}\Delta(W)^{n}_{m}).$
The objective becomes the minimization of the residuals such that $(R_{U})^{n}_{m}\rightarrow 0$. This gives us the criteria we need to form the following least squares problem:
$L_{U}=\mathbb{E}[||(R_{U})^{n}||_{2}^{2}].$ (22)
@@
$(v_{i,k}^{*})^{n}=\textnormal{arg}\max_{v\in\hat{\mathcal{V}}}\mathbb{E}_{v}[\hat{U}_{k}^{n}].$
@@
Given the optimal control $v_{i,k}^{*n}$ that maximizes (5), we can compute the expected gain and risk of the strategy by running a Monte Carlo simulation to compute the terminal wealth. The Monte Carlo follows the algorithm as shown in Tse [2012].
@@ 4 Neural network approach to the optimal trade execution problem
In this section, we propose a neural network approach to solve the system of least squares formulation (22). The method follows a general framework that has been used to solve high-dimensional American option pricing problems and has shown to produce good results Han and E [2016]; E et al. [2017]; Hure et al. [2020]; Chen and Wan [2021]; Na and Wan [2023]. This framework has also been used to solve more general HJB equations as shown in Hure et al. [2020].
@@ 4.1 Neural network formulation
We model the approximations $\hat{U}_{m}^{n}$ using a deep residual U-net with a self attention layer. U-net is a convolutional neural network architecture with great success in image processing and was first introduced for the application of biomedical image segmentation Ronneberger et al. [2015]. We chose this architecture to learn the relationship between the discretized trade rates and assets. This is because U-net has been shown to use and augment existing features more effectively Ronneberger et al. [2015]. We also explore the use of self-attention layer to increase the efficiency of the residual deep network by embedding temporal relationship between the previous estimate $(U)^{n+1}_{m}$ and the current states given by {$(S)^{n}_{m},(b)^{n}_{m},(\alpha)^{n}_{m}$, $(v)^{n}_{m}$}. The use of a self-attention layer allows our network to determine which features are the most relevant in predicting the output Vaswani et al. [2017]. This allows us to avoid using recurrent network structures which have a chance to exhibit vanishing gradient properties due to the activation functions used in their gate structures Vaswani et al. [2017].
@@ 4.1.1 Self-attention residual network
The standard attention network was initially used to model predictive sequential language models Vaswani et al. [2017]. They were initially implemented to learn the importance of features in predicting a language sequence. The self-attention architecture was introduced as a method to embed information inputs to create latent feature space for more efficient training Vaswani et al. [2017]. In this paper, we use self-attention layer to overcome the limitation of deep residual network methods. Deep residual networks need to save weights at every timestep Han and E [2016]; E et al. [2017]; Chen and Wan [2021]; however, by adding the self-attention layer allows us to capture the dependencies between features.
The attention mechanism relies on a query $Q$, and key-value pair $\{K,\tilde{V}\}$. We let $A(Q,K,\tilde{V})$ be the score function given by Vaswani et al. [2017]:
@@
where the softmax function is a standard activation function Goodfellow et al. [2016] and $l$ is the length of the sequence. Let $h=1,...,H$ be the number of heads where each head is given by $\mathcal{H}_{h}=A(Q,K,\tilde{V})$. It was found that constructing multiple attention layers in a bagging approach allowed for better performance than a single attention mechanism Vaswani et al. [2017]. The multi-head attention mechanism is the concatenation of $h$ heads and is defined as Vaswani et al. [2017]:
@@
where $\oplus$ is used to signify concatenation. It has been shown that using a linear layer to represent $\{Q,K,\tilde{V}\}$ is beneficial to learning Vaswani et al. [2017]. The self-attention network is a special case of the multi-head attention mechanism where the inputs are passed through a linear layer to construct $\{Q,K,\tilde{V}\}$. More specifically, given an input $X$ and weights $W_{Q}$, $W_{K}$, $W_{\tilde{V}}$ we can define $Q=W_{Q}X$, $K=W_{K}X$, $\tilde{V}=W_{\tilde{V}}X$.
The self-attention mechanism itself does not have much prediction power Vaswani et al. [2017]. To add prediction power to the self-attention layer, we couple it with a convolution neural network (CNN). The benefit of using a CNN is that it captures the high dimensional structure across the discretized space of trade rates and our input features Goodfellow et al. [2016]. More specifically, we use a U-net architecture coupled with a residual structure. The U-net we use has $4$ convolution layers and $4$ transposed convolution layers with a converging-diverging geometry as shown in Figure 1 Ronneberger et al. [2015]. The convergent-divergent structure of the U-net allows for the network to capture the content in the data and allows for localization Ronneberger et al. [2015]. It also allows us to limit the effects of features that may not contribute to predictions.
@@
In between each network layer, we normalize the input $X$. Normalization allows our network to learn more effectively and prevent cross correlation between network layers Goodfellow et al. [2016]. For learnable parameters $\varphi$ and $\eta$ to assist convergence of the deep residual network, we use group normalization, $G$, given by He et al. [2016]:
@@
For a given input, $X^{n}$, at timestep $n$, the residual U-net for each timestep has the following form He et al. [2016]:
@@
In our network, we use the swish activation function over the ReLU activation function. This is because the ReLU activation function may cause information loss when training over long sequences due to the negative values being zeroed out Na and Wan [2023]. The output of CNN plus some weighted $X_{a}$ gives us the final output:
@@ 4.2 Network Training and Prediction
Note, due to the non-uniqueness of the optimal control, we train multiple networks for each agent and perform bagging to ensure better results. This allows our neural network to explore multiple paths. Optimization is done using the Adam algorithm Kingma and Ba [2017] with a learning rate of $1\times 10^{-4}$.
@@ 4.3 Data Generation
where $v_{i,k}$ is in units of $1/T$. For $n=1,...,N$, the dynamics of (18) is simulated using Euler’s method. The dynamics of $\{(b_{k})^{n}_{m},(\alpha_{i,k})^{n}_{m}\}$ follow the dynamics given by (19) and (20). To ensure that all assets are liquidated at time $T$, we define the trade rate, $v_{T}$, at the instant before maturity as Forsyth [2011]:
@@ 5.1 Experiment 1: Comparison of Residual Attention Network to Existing solutions
In this experiment, we compare the approximation of the neural network solution to the results of existing work Forsyth [2011]; Aivaliotis and Veretennikov [2018]; Guan and Hu [2022]; Zhou and Li [2000]. This experiment verifies that the outputs of our neural network solution coincides with known solutions wherever they are available. We fix the parameter $\gamma=6,\phi=0.0$ and assume no price impact, i.e. $\kappa_{p}=0,\kappa_{\tau}=0$ and $\kappa_{s}=0$.
@@
In the case of a single agent with one asset, we compare the results to the FDM solution Forsyth [2011]; Aivaliotis and Palczewski [2014]. First, we look at the performance of the proposed method for different $N$. This is done to determine the appropriate number of timesteps we need for an accurate solution to the problem. We fix $M=3000$, $T=1$ year and run the neural network until $L_{U}(\Theta)<5\times 10^{-6}$ and summarize the results in Table 1. We also record the FDM solution for $N=10,100,200,300,400$ with relative error given by:
@@
Table 1: Residual U-net with self-attention approximation for the solution to the optimal value $u(x_{0},0)$ over different number of time-steps $N$. The FDM values are results from Aivaliotis and Palczewski [2014]. The relative errors are computed using (23)
@@
Table 2: Residual U-net with self-attention approximation for the solution to the mean-variance criterion $u(x_{0},0)$ over different number of simulations $M$. In the FDM solution $M$ represents the grid size used for the wealth Aivaliotis and Palczewski [2014]. The relative error is computed using (23)
We also make a comparison to the FDM solution to the optimal trade execution for $\kappa_{\tau}=2\times 10^{-6}$, $T=1/250$, $\mu=0$ and $r=0$ Forsyth [2011]. We compute the efficient frontier Tse [2012] as shown in Figure 3. For an initial price of $s(0)=100$ and $\alpha(0)=1.0$, the expected gain from selling is $99.295$ with a standard deviation of $0.7469$ Forsyth [2011]. Our neural network approximation gives an expected gain of $99.271$ with a standard deviation of $0.7402$, similar to the FDM results.
Figure 3: The efficient frontier interpolated from the neural network approximation (blue) and interpolated values from Forsyth [2011] (orange). Interpolation was performed using a polynomial degree $4$.
@@
We compare the neural network solution with analytical solutions. For the special case of no price impact, an analytical solution exists for the case of a single agent with multiple assets Zhou and Li [2000]. This experiment shows that our outputs are consistent with multidimensional results.
Let $A=[\mu_{1}-r,....,\mu_{d}-r]^{\top},\underline{B}=[\sigma_{1},...,\sigma_{d}]^{\top}$ be the vector of drift rates and the covariance matrix for each assets $i=1,...,d$. Note this framework assumes independence between the assets, and hence we represent the volatilities as a vector. The time-consistent mean variance optimal asset allocation for multiple assets is given by Zhou and Li [2000]:
@@
where parameter $\mu>0$ and $-\infty<\lambda<\infty$. Let, $\mu=1$, $\lambda=\gamma$ and let $\alpha^{*}(t)$ be the quantity of assets held by the agent. Then the optimal asset allocation for a single agent with multiple assets is given by Zhou and Li [2000]:
@@
where $\underline{X}(t)=[X_{1}(t),...,X_{d}(t)]^{\top}.$ Our model can be used to compute $\alpha^{*}(t)$ by solving the equation $\alpha^{*}(t)=\alpha^{*}(t-dt)+v^{*}(t)dt$. We fix $N=100$, and $M=1000$ and compute $\mathbb{E}[J(X(T));\alpha^{*}(t),t=0,\Delta t,...,N\Delta t]$ using Monte Carlo simulation Zhou and Li [2000]. Given $\{\alpha_{1}(0)=1.0,\alpha_{2}(0)=0.5\}$, the optimal value for $d=2$ at $T=1/250,1/52,1/12,1$ years are summarized in Table 3. We see that the highest relative error occurs at $T=1$ with a relative error of $0.138\%$.
@@
We compare the neural network solution with analytical solutions. For the special case of no price impact, an analytical solution exists for multiple agent with one asset Guan and Hu [2022]. The mean-variance optimal portfolio problem can be written as Guan and Hu [2022]:
@@
The analytical solution to the time-consistent auxiliary problem is given by Guan and Hu [2022]:
@@
We use Monte Carlo Simulation to compute $\mathbb{E}[J(X(T));\alpha^{*}(t),t=0,\Delta t,...,N\Delta t]$ Guan and Hu [2022]. Given $\alpha_{k}(0)=1.0$ the optimal value for $K=2$ are summarized in Table 4. We present results for only one agent as the agents are assumed to be symmetric. We can see from Table 4 that the relative errors are generally small with the highest relative error of $0.0351\%$ occurred at $T=1/12$.
@@ 5.2 Experiment 2: Sensitivity of Parameters
To analyze the sensitivity of $\phi$, we let $\gamma=6$, $\kappa_{p}=0$ and $\kappa_{\tau}=0$. The results are shown in Figure 4. As we can see, as $\phi$ increases, $\alpha^{*}$ also increases. This is expected since the agent invests more into risky assets as they become more aware of other agents Guan and Hu [2022]. Thus our neural network approach is able to capture this behaviour correctly.
@@
Moreover, we also want to ensure that $\alpha^{*}$ decreases as $\gamma$ increases. This behaviour is an important behaviour to observe in agents as it is commonly known that as the risk aversion increases, the agent invests less in risky assets Zhou and Li [2000]. We fix $\phi=0.0$, $\kappa_{p}=0.0$, and $\kappa_{\tau}=0.0$. Figure 5 shows that $\alpha^{*}$ indeed decreases as $\gamma$ increases, illustrating the effectiveness of the neural network approach.

### 19 Mar 2024 v2

@@ 3.3.1 Least squares formulation
$(R_{U})^{n}_{m}=(U_{k})^{n+1}_{m}-((\hat{U}_{k})^{n}_{m}-\left(D_{t}(\hat{U}_{k})^{n}_{m}+(\mathcal{L}\hat{U}_{k})^{n}_{m}\right)\Delta t+\sum\limits_{i=1}^{d}\sigma_{i}D_{S_{i}}(\hat{U}_{k})^{n}_{m}\Delta(W)^{n}_{m}).$
The objective becomes the minimization of the residuals such that $(R_{U})^{n}_{m}\rightarrow 0$. Since our $K$ BSDE equations are coupled by the optimal control and solved backwards in time, we satisfy the optimality condition which solves the multi-agent game as in Definition 2.1.
This gives us the criteria we need to form the following least squares problem:
$L_{MSE}=\mathbb{E}[||(R_{U})^{n}||_{2}^{2}].$ (20)
@@
$(v_{i,k}^{*})^{n}=\textnormal{arg}\max_{v\in\hat{\mathcal{V}}}\mathbb{E}_{v}[(\hat{U}_{k})^{n}].$
@@
Given the optimal control $v_{i,k}^{*n}$ that maximizes (5), we can compute the expected gain and risk of the strategy by running a Monte Carlo simulation to compute the terminal wealth. The Monte Carlo follows the algorithm as shown in [Tse 2012].
@@ 4 Neural network approach to the optimal trade execution problem
In this section, we propose a neural network approach to solve the system of least squares formulation (20). The method follows a general framework that has been used to solve high-dimensional American option pricing problems and has shown to produce good results [Han and E 2016; E et al. 2017; Hure et al. 2020; Chen and Wan 2021; Na and Wan 2023]. This framework has also been used to solve more general HJB equations as shown in [Hure et al. 2020].
@@ 4.1 Neural network formulation
We model the approximations $\hat{U}_{m}^{n}$ using a deep residual U-net with a self attention layer. U-net is a convolutional neural network architecture with great success in image processing and was first introduced for the application of biomedical image segmentation [Ronneberger et al. 2015]. We chose this architecture to learn the relationship between the discretized trade rates and assets. This is because U-net has been shown to use and augment existing features more effectively [Ronneberger et al. 2015]. We also explore the use of self-attention layer to increase the efficiency of the residual deep network by embedding temporal relationship between the previous estimate $(U)^{n+1}_{m}$ and the current states given by {$(S)^{n}_{m},(b)^{n}_{m},(\alpha)^{n}_{m}$, $(v)^{n}_{m}$}. The use of a self-attention layer allows our network to determine which features are the most relevant in predicting the output [Vaswani et al. 2017]. This allows us to avoid using recurrent network structures which have a chance to exhibit vanishing gradient properties due to the activation functions used in their gate structures [Vaswani et al. 2017].
@@ 4.1.1 Self-attention residual network
The standard attention network was initially used to model predictive sequential language models [Vaswani et al. 2017]. They were initially implemented to learn the importance of features in predicting a language sequence. The self-attention architecture was introduced as a method to embed information inputs to create latent feature space for more efficient training [Vaswani et al. 2017]. In this paper, we use self-attention layer to overcome the limitation of deep residual network methods. Deep residual networks need to save weights at every timestep [Han and E 2016; E et al. 2017; Chen and Wan 2021]; however, by adding the self-attention layer allows us to capture the dependencies between features.
The attention mechanism relies on a query $Q$, and key-value pair $\{K,\tilde{V}\}$. We let $A(Q,K,\tilde{V})$ be the score function given by [Vaswani et al. 2017]:
@@
where the softmax function is a standard activation function [Goodfellow et al. 2016] and $l$ is the length of the sequence. Let $h=1,...,H$ be the number of heads where each head is given by $\mathcal{H}_{h}=A(Q,K,\tilde{V})$. It was found that constructing multiple attention layers in a bagging approach allowed for better performance than a single attention mechanism [Vaswani et al. 2017]. The multi-head attention mechanism is the concatenation of $h$ heads and is defined as [Vaswani et al. 2017]:
@@
where $\oplus$ is used to signify concatenation. It has been shown that using a linear layer to represent $\{Q,K,\tilde{V}\}$ is beneficial to learning [Vaswani et al. 2017]. The self-attention network is a special case of the multi-head attention mechanism where the inputs are passed through a linear layer to construct $\{Q,K,\tilde{V}\}$. More specifically, given an input $X$ and weights $W_{Q}$, $W_{K}$, $W_{\tilde{V}}$ we can define $Q=W_{Q}X$, $K=W_{K}X$, $\tilde{V}=W_{\tilde{V}}X$.
The self-attention mechanism itself does not have much prediction power [Vaswani et al. 2017]. To add prediction power to the self-attention layer, we couple it with a convolution neural network (CNN). The benefit of using a CNN is that it captures the high dimensional structure across the discretized space of trade rates and our input features [Goodfellow et al. 2016]. More specifically, we use a U-net architecture coupled with a residual structure. The U-net we use has $4$ convolution layers and $4$ transposed convolution layers with a converging-diverging geometry as shown in Figure 1 [Ronneberger et al. 2015]. The convergent-divergent structure of the U-net allows for the network to capture the content in the data and allows for localization [Ronneberger et al. 2015]. It also allows us to limit the effects of features that may not contribute to predictions.
@@
In between each network layer, we normalize the input $X$. Normalization allows our network to learn more effectively and prevent cross correlation between network layers [Goodfellow et al. 2016]. For learnable parameters $\varphi$ and $\eta$ to assist convergence of the deep residual network, we use group normalization, $G$, given by [He et al. 2016]:
@@
For a given input, $X^{n}$, at timestep $n$, the residual U-net for each timestep has the following form [He et al. 2016]:
@@
In our network, we use the swish activation function over the ReLU activation function. This is because the ReLU activation function may cause information loss when training over long sequences due to the negative values being zeroed out [Na and Wan 2023]. The output of CNN plus some weighted $X_{a}$ gives us the final output:
@@ 4.2 Network Training and Prediction
Note, due to the non-uniqueness of the optimal control, we train multiple networks for each agent and perform bagging to ensure better results. This allows our neural network to explore multiple paths. Optimization is done using the Adam algorithm [Kingma and Ba 2017] with a learning rate of $1\times 10^{-4}$.
@@ 4.3 Data Generation
where $v_{i,k}$ is in units of $1/T$. For $n=1,...,N$, the dynamics of (16) is simulated using Euler’s method. The dynamics of $\{(b_{k})^{n}_{m},(\alpha_{i,k})^{n}_{m}\}$ follow the dynamics given by (17) and (18). To ensure that all assets are liquidated at time $T$, we define the trade rate, $v_{T}$, at the instant before maturity as [Forsyth 2011]:
@@ 5.1 Experiment 1: Comparison of Residual Attention Network to Existing solutions
In this experiment, we compare the approximation of the neural network solution to the results of existing work [Forsyth 2011; Aivaliotis and Veretennikov 2018; Guan and Hu 2022; Zhou and Li 2000]. This experiment verifies that the outputs of our neural network solution coincides with known solutions wherever they are available. We fix the parameter $\gamma=6,\phi=0.0$ and assume no price impact, i.e. $\kappa_{p}=0,\kappa_{\tau}=0$ and $\kappa_{s}=0$.
@@
In the case of a single agent with one asset, we compare the results to the FDM solution [Forsyth 2011; Aivaliotis and Palczewski 2014]. First, we look at the performance of the proposed method for different $N$. This is done to determine the appropriate number of timesteps we need for an accurate solution to the problem. We fix $M=3000$, $T=1$ year and run the neural network until $L_{U}(\Theta)<5\times 10^{-6}$ and summarize the results in Table 1. We also record the FDM solution for $N=10,100,200,300,400$ with relative error given by:
@@
Table 1: Residual U-net with self-attention approximation for the solution to the optimal value $u(x_{0},0)$ over different number of time-steps $N$. The FDM values are results from [Aivaliotis and Palczewski 2014]. The relative errors are computed using (21)
@@
Table 2: Residual U-net with self-attention approximation for the solution to the mean-variance criterion $u(x_{0},0)$ over different number of simulations $M$. In the FDM solution $M$ represents the grid size used for the wealth [Aivaliotis and Palczewski 2014]. The relative error is computed using (21)
We also make a comparison to the FDM solution to the optimal trade execution for $\kappa_{\tau}=2\times 10^{-6}$, $T=1/250$, $\mu=0$ and $r=0$ [Forsyth 2011]. We compute the efficient frontier [Tse 2012] as shown in Figure 3. For an initial price of $s(0)=100$ and $\alpha(0)=1.0$, the expected gain from selling is $99.295$ with a standard deviation of $0.7469$ [Forsyth 2011]. Our neural network approximation gives an expected gain of $99.271$ with a standard deviation of $0.7402$, similar to the FDM results.
Figure 3: The efficient frontier interpolated from the neural network approximation (blue) and interpolated values from [Forsyth 2011] (orange). Interpolation was performed using a polynomial degree $4$.
@@
We compare the neural network solution with analytical solutions. For the special case of no price impact, an analytical solution exists for the case of a single agent with multiple assets [Zhou and Li 2000]. This experiment shows that our outputs are consistent with multidimensional results.
Let $A=[\mu_{1}-r,....,\mu_{d}-r]^{\top},\underline{B}=[\sigma_{1},...,\sigma_{d}]^{\top}$ be the vector of drift rates and the covariance matrix for each assets $i=1,...,d$. Note this framework assumes independence between the assets, and hence we represent the volatilities as a vector. The time-consistent mean variance optimal asset allocation for multiple assets is given by [Zhou and Li 2000]:
@@
where parameter $\mu>0$ and $-\infty<\lambda<\infty$. Let, $\mu=1$, $\lambda=\gamma$ and let $\alpha^{*}(t)$ be the quantity of assets held by the agent. Then the optimal asset allocation for a single agent with multiple assets is given by [Zhou and Li 2000]:
@@
where $\underline{X}(t)=[X_{1}(t),...,X_{d}(t)]^{\top}.$ Our model can be used to compute $\alpha^{*}(t)$ by solving the equation $\alpha^{*}(t)=\alpha^{*}(t-dt)+v^{*}(t)dt$. We fix $N=100$, and $M=1000$ and compute $\mathbb{E}[J(X(T));\alpha^{*}(t),t=0,\Delta t,...,N\Delta t]$ using Monte Carlo simulation [Zhou and Li 2000]. Given $\{\alpha_{1}(0)=1.0,\alpha_{2}(0)=0.5\}$, the optimal value for $d=2$ at $T=1/250,1/52,1/12,1$ years are summarized in Table 3. We see that the highest relative error occurs at $T=1$ with a relative error of $0.138\%$.
@@
We compare the neural network solution with analytical solutions. For the special case of no price impact, an analytical solution exists for multiple agent with one asset [Guan and Hu 2022]. The mean-variance optimal portfolio problem can be written as [Guan and Hu 2022]:
@@
The analytical solution to the time-consistent auxiliary problem is given by [Guan and Hu 2022]:
@@
We use Monte Carlo Simulation to compute $\mathbb{E}[J(X(T));\alpha^{*}(t),t=0,\Delta t,...,N\Delta t]$ [Guan and Hu 2022]. Given $\alpha_{k}(0)=1.0$ the optimal value for $K=2$ are summarized in Table 4. We present results for only one agent as the agents are assumed to be symmetric. We can see from Table 4 that the relative errors are generally small with the highest relative error of $0.0351\%$ occurred at $T=1/12$.
@@ 5.2 Experiment 2: Sensitivity of Parameters
To analyze the sensitivity of $\phi$, we let $\gamma=6$, $\kappa_{p}=0$ and $\kappa_{\tau}=0$. The results are shown in Figure 4. As we can see, as $\phi$ increases, $\alpha^{*}$ also increases. This is expected since the agent invests more into risky assets as they become more aware of other agents [Guan and Hu 2022]. Thus our neural network approach is able to capture this behaviour correctly.
@@
Moreover, we also want to ensure that $\alpha^{*}$ decreases as $\gamma$ increases. This behaviour is an important behaviour to observe in agents as it is commonly known that as the risk aversion increases, the agent invests less in risky assets [Zhou and Li 2000]. We fix $\phi=0.0$, $\kappa_{p}=0.0$, and $\kappa_{\tau}=0.0$. Figure 5 shows that $\alpha^{*}$ indeed decreases as $\gamma$ increases, illustrating the effectiveness of the neural network approach.